a ball is thrown upward from the ground at an angle of 60° with the horizontal with a velocity of 6m/s. a)find its vertical postion with respect to the ground when it has travelled a distance of 2m horizontally. b) find its velocity(upward or downward) at this point


1
Expert's answer
2021-09-07T09:50:24-0400

Given:

"\theta=60^{\circ}"

"v_0=6\:\rm m/s"

"x=2\:\rm m"


(a) the equation of trajectory of a ball is given by

"y=x\tan\theta-\frac{gx^2}{2v_0^2\cos^2\theta}"

So

"y=2\tan 60^{\circ}-\frac{9.8*2^2}{2*6^2*\cos^260^{\circ}}=1.3\:\rm m"

(b)

"x=v_0\cos\theta*t"

"t=\frac{x}{v_0\cos\theta}=\frac{2}{6*\cos60^{\circ}}=0.67\:\rm s"

The components of velocity vector are

"v_x=v_0\cos\theta=6*\cos60^{\circ}=3\:\rm m/s"

"v_y=v_0\sin\theta-gt=6*\sin60^{\circ}-9.8*0.67\\
=-1.3\:\rm m/s"

The velocity is directed downward.